OpenStudy (anonymous):

A stone is dropped into a lake, creating a circular ripple that travels outward at a speed 25cm/sec. Find the rate at which the area within the circle is increasing after 4sec.

OpenStudy (ybarrap):

$$ A=\pi r^2\\ {dA\over dt}=2\pi r{dr\over dt}=2\pi rv\\ r=\int v~dt=vt+K\\ r(0)=0\implies K=0\\ r(t)=vt\\ \implies {dA\over dt}=2\pi r{dr\over dt}=2\pi v^2t\\ $$ \(v=0.25\) and \(t=4\) Does this make sense?

OpenStudy (anonymous):

sorry, no. I'm doing basic related rates...

OpenStudy (ybarrap):

Do you know calculus?

OpenStudy (anonymous):

Well, i just started taking the class. I'm on the application of derivatives.

OpenStudy (ybarrap):

Ok, so first we define Area Next we define its derivative with respect to time $$ A=\pi r^2\\ {dA\over dt}=2\pi r{dr\over dt}\\ $$ The only thing that changes is \(r\), so that is the only thing about Area that can change. Does this make sense?

OpenStudy (anonymous):

Yes

OpenStudy (ybarrap):

So since \(r\) is the only thing that can change and we also know that \(r\) is a function of time; when we take the derivative of A with respect to time, we also take the derivative of \(r\) with respect to time, because that is the thing about Area that is changing. The derivative of \(r^2\) is \(2r\). But \(r\) is also a function of time, so we must multiply \(2r\) by \({dr\over dt}\) to account for this fact. $$ {dA\over dt}=2\pi r{dr\over dt}\\ $$ We also know that \({dr\over dt}\) is just velocity (distance traveled over time taken). So \({dr\over dt}=v\). Ok so far?

OpenStudy (anonymous):

i think so... okay.

OpenStudy (ybarrap):

Note that we used a rule called "Implicit Differentiation" when we took the derivative of A: $$ {dA\over dt}=2\pi r{dr\over dt}\\ $$ http://en.wikipedia.org/wiki/Implicit_function#Implicit_differentiation It might be better to write \(A\) and \(r\) as a function of time \(A=A(t),r=r(t)\) to remind us that we will need to use this rule: $$ {dA(t)\over dt}=2\pi r(t){dr(t)\over dt}\\ $$

OpenStudy (ybarrap):

Does velocity make sense to you? Do you see that \({dr\over dt}\) is just velocity?

OpenStudy (anonymous):

I think so, because the radius represents the distance, so there is distance over time, which the velocity.

OpenStudy (ybarrap):

Next we need to find a relationship between \(r\) and time. To do this, we know that velocity times time is distance (using the same idea that velocity is distance over time). Therefore, $$ r(t)=vt $$ This is how big the circle is after some time \(t\). We now have everything we need. $$ {dA\over dt}=2\pi r{dr\over dt}=2\pi \times vt\times v\\ $$

OpenStudy (ybarrap):

This represents how fast the Area \(A\) is changing \(dA\) over any change in time \(dt\).

OpenStudy (ybarrap):

It takes time to process all this material. Read over it and ask any questions you want later. I'll check over the next few days. Ok?

OpenStudy (anonymous):

okay, thank you!!

OpenStudy (ybarrap):

You're welcome. Keep hacking at this, took me some time as well, but you will get it!!

OpenStudy (anonymous):

dA/dt=2pi(100)(25)??? @ybarrap

OpenStudy (ybarrap):

0.25, not 25, answer will be in meters squared per second

OpenStudy (anonymous):

But all the values are in centimeters. @ybarrap